OpenStudy (anonymous):

Literal equation please help step by step!The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the force of gravity, and t = time. Solve this equation for g, and use your formula to determine the force of gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 feet. Show all steps in your work.

OpenStudy (anonymous):

@Hero

hero (hero):

d = 490 g = 10 490 = 0.5g(10)^2 Finish solving for g

hero (hero):

@ris_lynn

OpenStudy (anonymous):

how do I solve for g :O and can ushow me how you got that! please @Hero

hero (hero):

Well, they mentioned that the ball was dropped from 490 ft and they said it will hit the ground. The ground represents a height of zero. The distance from the ground to the height where the ball will be dropped from is 490 ft. So d = 490. Next they said it will take a time of 10 seconds for the ball to hit the ground so t = 10. So you plug those values in to the formula d = 0.5gt^2, then solve for g

hero (hero):

To solve for g: 490 = 0.5g(10)^2 490 = (0.5)(100)g 490 = 50g 490/50 = g 9.8 = g

OpenStudy (anonymous):

oh okay is that all I do ?

hero (hero):

Yes

OpenStudy (anonymous):

thankyouuu!